XYZ Inc. has a machine with a book value of $50,000 and a five-year remaining life....

  1. XYZ Inc. has a machine with a book value of $50,000 and a five-year remaining life. A new machine is available at a cost of $85,000 and XYZ can also receive $38,000 for trading in the old machine. The new machine will reduce variable manufacturing costs by $14,000 per year over its five-year life. Should the machine be replaced?
Answer #1

Computation of Incremental cost.

Cost of New Machine = ($85,000)

salvage value of old machine = $38,000

Net incremental cost = ($47,000)

Net benefits due to reduction in variable Manufacturing costs = $14,000 * 5 = $70,000.

So, Income will increase by = 70,000 - 47,000 = $23,000

Yes, the machine should be replaced. because Income will increase by $23,000 in total